OpenStudy (anonymous):

A prism has a right triangle as its base with a hypotenuse that measures 10 meters and legs that measure 6 meters and 8 meters. What is the surface area of this prism if the height of the prism is 8 meters?

hero (hero):

In order to solve this, you'll need to know 1. The Surface Area of a Prism Formula 2. How to find the area of the base 3. How to apply the dimensions to the Prism Formula 4. How to solve equations Which of the above are you having trouble with?

OpenStudy (anonymous):

How to solve equations

hero (hero):

So basically, what you mean to tell me is, 1. You've already written down the surface area of the the prism 2. You've already found the area of the base 3. And you've already applied the dimensions of the figure to the formula and now all you need to do is solve the equation, correct?

Directrix (directrix):

The surface area of prism is the perimeter of its base times its height + two times the area of its base. SA = ph + 2B SA = (6 + 8 + 10)*8 + 2 (1/2)(6)(8) (Note that the base is a right triangle with area one-half the product of its legs of 6 and 8.) SA = 24*8 + 48 SA = 192 + 48 = 240 square meters